# PALADIN : Robust Neural Fingerprinting for Text-to-Image Diffusion Models

### v1

@@ Abstract
The risk of misusing text-to-image generative models for malicious uses, especially due to the open-source development of such models, has become a serious concern. As a risk mitigation strategy, attributing generative models with neural fingerprinting is emerging as a popular technique. There has been a plethora of recent work that aim for addressing neural fingerprinting. A trade-off between the attribution accuracy and generation quality of such models has been studied extensively. None of the existing methods yet achieved $100\%$ attribution accuracy. However, any model with less than perfect accuracy is practically non-deployable. In this work, we propose an accurate method to incorporate neural fingerprinting for text-to-image diffusion models leveraging the concepts of cyclic error correcting codes from the literature of coding theory.
@@ 1 Introduction
The current state-of-the-art text-to-image generative models have shown prodigious capabilities in generating high-quality and diverse images from natural language text prompts. This has enabled the widespread creation and manipulation of photo realistic images, adhering to precise textual prompts. This has led to many image editing tools [2] [4] [11] that are becoming creative resources for artists, designers, and the public. Although this is a great leap forward for generative modeling, it raises concerns about the authenticity, veracity, validity, and provenance of the generated images [5]. The increasing realism of such generated images poses a serious threat to both individual and social interests. A convincing and widely shared misinformation could disrupt democratic discussion, influence elections, or threaten national security. A viable solution to this problem is to create accountability for the generated image by integrating user information in the generated artifacts.
A line of work in this area has emerged that analyze design principles of fingerprinting strategies in text-to-image diffusion models and the trade-off between the attribution accuracy and generated image quality. However, none of such methods is able to achieve 100% accuracy. For example, an attribution accuracy of 0.99% means the aforesaid method will still predict $100,000$ user incorrectly for a database containing 1 billion users. Unless a method produces absolute zero incorrect predictions, such near-perfect methods can not be deployed in the real world.
Figure 1: PALADIN: a robust and model-agnostic strategy that elevates a near-perfect latent diffusion based fingerprinting strategy to 100% attribution accuracy utilizing binary cyclic error correction codes.
@@
  • We propose a framework, PALADIN, for accurate fingerprinting to elevate a near-perfect model to a perfect one utilizing cyclic error correction codes.
@@
To the best of our knowledge, this is the first method that achieves 100% attribution accuracy and the lowest FER for the text-to-image diffusion model.
@@ 2 Related Work
One of the recent state-of-the-art method ,WOUAF [7], introduced the concept of embedding the watermarks in all the images they generate by fine-tuning of the decoder of Latent Diffusion Models. WOUAF [7] is based on StyleGAN [6] weight modulation technique to embed user information in the generated artifacts. These watermarks are more robust than previously existing models could generate, invisible to the human eye and could be employed to detect generated images and identify the user that generated it, with high performance. In ours, we make use of recent progress in fingerprinting in diffusion models and enhance their effectiveness with the power of error correction capability. PALADIN can broadly be classified as distributor-oriented fingerprinting, in which the model distributor assigns and embeds a unique traceable fingerprint to each user who downloads the model. In particular, we build atop WOUAF. However, the core principle can be applied to other models as well.
@@ 3.4 Fingerprint decoder
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG Comp.
(1.35) (1.35) (1.35) (1.35) $(\sigma=0.1)$ (20%) (qual. = 60)
WOUAF 0.99872 0.99884 0.99887 0.99281 0.99744 0.99910 0.99754 0.99004
PALADIN 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000
(a) Table shows the fingerprint accuracy against different image post processing.
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG Comp.
(1.35) (1.35) (1.35) (1.35) $(\sigma=0.1)$ (20%) (qual. = 60)
WOUAF* 0.03622 0.03302 0.03342 0.08421 0.07045 0.02301 0.06745 0.22357
PALADIN 0.00080 0.00060 0.00100 0.00800 0.00060 0.00120 0.01120 0.01241
(b) Table shows the fingerprint error rate against different image post processing. *WOUAF does not report fingerprint corruption, all the data calculated are on testbench with reference fingerprint.
Table 4: Accuracy and FER readings against varied post-processing operations.
@@ 3.5 Loss function
Ciphering network $\mathcal{E_{\psi}}$, SD decoder $\mathcal{D}$ and the deciphering network $\mathcal{D_{\psi}}$ are jointly optimized to ensure robust fingerprinting.
@@ 4.2 Evaluation
As demonstrated in Table 1, PALADIN outperforms WOUAF (our strongest baseline) significantly and produces robust and superior performance as per all metrics. Even pre-error correction, PALADIN $\psi$’s bit accuracy is significantly higher than WOUAF. A key feature of PALADIN is the ability to detect and report fingerprint corruption.
@@
As shown in Table 2, the quality of generated images with PALADIN is closer to the original SD’s compared to WOUAF as measured by by SSIM, PSNR, LPIPS, and FID. Table 3 shows image quality assessment for PALADIN compared to Stable Diffusion and WOUAF. PALADIN performs better than WOUAF in terms of SSIM and PSNR, while improves upon SD in terms of LPIPS and FID. Figure 2 shows qualitative results which highlight PALADIN’s ability to reliably incorporate fingerprinting without degrading image generation quality.
[Plot showing bit accurary for different brightness scale] [Plot showing FER for different brightness scale]
[Plot showing bit accurary for different contrast scale] [Plot showing FER for different contrast scale]
[Plot showing bit accurary for different sharpness scale] [Plot showing FER for different sharpness scale]
[Plot showing bit accurary for different saturation scale] [Plot showing FER for different saturation scale]
Figure 3: Plots showing accuracy and FER for varying degree of post processing on images.
[Plot showing bit accurary for different jpeg compression scale] [Plot showing FER for different jpeg compression scale]
Figure 4: Plots showing accuracy and FER for varying degree of jpeg compression on images.
@@ 4.3 Resilience against image post-processing
One of the most common techniques to corrupt or erase fingerprinting is image post-processing techniques, in that a malicious user attempts to obscure fingerprint by employing one or more post processing methods on generated image whilst preserving the photo-realistic nature of image. Here, we test PALADIN against multiple post processing attacks such as brightness, contrast, saturation, sharpness, horizontal flip, crop and gaussian noise and benchmark against WOUAF. Table 4(a) and 4(b) demonstrates PALADIN robust accuracy and FER respectively against WOUAF on all image post processing employed. PALADIN clearly shows stable accuracy while maintaining low FER. Figure 3 shows accuracy variation versus the strength of the post processing, PALADIN shows stable performance and maintains constant bit accuracy and FER over a range of post processing operations normally used to erase or corrupt fingerprint.
  We also evaluate robustness against JPEG compression, from Table 4(a) and 4(b) we infer that PALADIN outperforms WOUAF when employed with JPEG compression with jpeg quality set to 60. Figure 4 shows accuracy variation versus the jpeg compression strenght, PALADIN shows stable performance and maintains constant bit accuracy and FER over a range of jpeg compression, thus proving if incorporated during training PALADIN generalizes well and outperforms WOUAF.
@@ 5 Conclusions
Near-perfect fingerprinting strategies for latent diffusion models, even with 99% attribution accuracies, cannot be deployed due to the unacceptably large absolute number of incorrect predictions involved. We introduce a framework, PALADIN, which preserves generated image quality and also achieves 100% user-attribution accuracy. PALADIN is developed on top of a recent state-of-the-art model, WOUAF. We show how the performance of WOUAF can be further improved by a combination of modified fingerprint decoder architecture and a loss function better equipped to handle the trade-off between generated image quality and the attribution accuracy. Next, we further enhance the model performance by incorporating the ability of bit error recovery, a concept borrowed from the literature of coding theory. Although PALADIN is built atop WOUAF, it is not restricted to this one model. PALADIN provides a plug-and-play layer of cyclic error correction scheme on top of similar models and is capable of making them 100% user-attribution accurate, along with an ability to appropriately flag non-identifiable fingerprints. In this work, we also introduce a new metric FER for measuring fingerprinting accuracy, that is more accurate and informative. To the best of our knowledge, PALADIN is the first ever strategy to achieve 100% user attribution accuracy in text-to-image diffusion models, which makes it possible for real-world deployment.

### v2

@@ Abstract
The risk of misusing text-to-image generative models for malicious uses, especially due to the open-source development of such models, has become a serious concern. As a risk mitigation strategy, attributing generative models with neural fingerprinting is emerging as a popular technique. There has been a plethora of recent work that aim for addressing neural fingerprinting. A trade-off between the attribution accuracy and generation quality of such models has been studied extensively. None of the existing methods yet achieved $100\%$ attribution accuracy. However, any model with less than cent percent accuracy is practically non-deployable. In this work, we propose an accurate method to incorporate neural fingerprinting for text-to-image diffusion models leveraging the concepts of cyclic error correcting codes from the literature of coding theory.
@@ 1 Introduction
The current state-of-the-art text-to-image generative models have shown significant capabilities in generating high-quality and diverse images from natural language text prompts. This has enabled the widespread creation and manipulation of photo realistic images, adhering to precise textual prompts. This has led to many image editing tools [2] [4] [13] that are becoming creative resources for artists, designers, and the public. Although this is a great leap forward for generative modeling, it raises concerns about the authenticity, veracity, validity, and provenance of the generated images [5]. The increasing realism of such generated images poses a serious threat to both individual and social interests. A convincing and widely shared misinformation could disrupt democratic discussion, influence elections, or threaten national security. A viable solution to this problem is to create accountability for the generated image by integrating user information in the generated artifacts.
A line of work in this area has emerged that analyze design principles of fingerprinting strategies in text-to-image diffusion models and the trade-off between the attribution accuracy and generated image quality. However, none of such methods is able to achieve 100% accuracy. For example, an attribution accuracy of 99.9% means the aforesaid method will still predict $1\ million$ user incorrectly for a database containing $1\ billion$ users, such methods can not be deployed in the real world.
Figure 1: PALADIN training pipeline. The weights of the denoising unet and text encoder are frozen and not part of training, weights of Cipher network, decoder part of Stable Diffusion and decipher network are updated as part of training. The attack block performs image post-processing such as Brightness, Gaussian Noise, H Flip, and others randomly with a probability of 50%. BCH encoder and decoder block is configured to codeword length = 63, message length = 39, capable of detecting and correcting up to 4-bit errors.
@@
  • We propose a framework, PALADIN, for accurate fingerprinting utilizing cyclic error correction codes.
@@ 2 Related Work
One of the recent state-of-the-art method ,WOUAF [7], introduced the concept of embedding the watermarks in all the images they generate by fine-tuning of the decoder of Latent Diffusion Models. WOUAF [7] is based on StyleGAN [6] weight modulation technique to embed user information in the generated artifacts. These watermarks are more robust than previously existing models could generate, invisible to the human eye and could be employed to detect generated images and identify the user that generated it, with high performance. In ours, we make use of recent progress in fingerprinting in diffusion models and enhance their effectiveness with the power of error correction capability. PALADIN can broadly be classified as distributor-oriented fingerprinting, in which the model distributor assigns and embeds a unique traceable fingerprint to each user who downloads the model. In particular, we build atop WOUAF [7]. However, the core principle can be applied to other models as well.
@@ 3.4 Fingerprint decoder
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG
[0.7,1.3] [0.7,1.3] [0.7,1.3] [0.7,1.3] $\sigma\in$[0,0.1] [0,0.2] $\in$[60,99]
WOUAF 0.9970 0.9969 0.9972 0.9928 0.9944 0.9991 0.9975 0.9900
PALADIN 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000 1.0000
(a) The table shows the minimum fingerprint accuracy seen against different image post-processing for varied strength.
Method Brightness Contrast Saturation Sharpness H Flip G Noise Crop JPEG
[0.7,1.3] [0.7,1.3] [0.7,1.3] [0.7,1.3] $\sigma\in$[0,0.1] [0,0.2] $\in$[60,99]
WOUAF 0.0362 0.0330 0.0334 0.0842 0.0704 0.0230 0.0674 0.2235
PALADIN 0.0008 0.0006 0.0010 0.0080 0.0006 0.0012 0.0112 0.0124
(b) The table shows the maximum fingerprint error rate seen against different image post-processing for varied strength.
Table 4: Accuracy and FER readings against varied image post-processing operations. Note accuracy drop is seen beyond above mentioned post-processing scale values, but image quality also gets affected, implying fingerprint corruption at the cost of reduced image quality.
@@ 3.5 Loss function
The ciphering network $\mathcal{E_{\psi}}$, SD decoder $\mathcal{D}$ and the deciphering network $\mathcal{D_{\psi}}$ are jointly optimized to ensure robust fingerprinting.
@@ 4.2 Evaluation
As demonstrated in Table 1, PALADIN outperforms WOUAF (our strongest baseline) significantly and produces robust and superior performance as per all metrics. Even pre-error correction, PALADIN $\psi$’s bit accuracy is significantly higher than WOUAF. A key feature of our methodology is the ability to detect and report fingerprint corruption.
@@
As shown in Table 2, the quality of generated images with our methodology is closer to the original SD’s compared to WOUAF as measured by SSIM, PSNR, LPIPS, and FID. Table 3 shows image quality assessment for our methodology compared to Stable Diffusion and WOUAF. PALADIN performs better than WOUAF in terms of SSIM and PSNR, while improves upon SD in terms of LPIPS and FID. Figure 2 shows qualitative results that highlight PALADIN’s ability to reliably incorporate fingerprinting without degrading image generation quality.
DDIM Euler
T = 40 T = 50 T = 60 T = 15 T = 20 T = 25
WOUAF 0.99243 0.99644 0.99036 0.99382 0.99420 0.99007
PALADIN 1.00000 1.00000 1.00000 1.00000 1.00000 1.00000
(a) Table showing effect of different scheduler on fingerprint accuracy.
DDIM Euler
T = 40 T = 50 T = 60 T = 15 T = 20 T = 25
WOUAF 0.17137 0.17741 0.20967 0.16330 0.07002 0.15725
PALADIN 0.00980 0.00820 0.01120 0.01020 0.00720 0.01120
(b) Table showing effect of different scheduler on fingerprint accuracy.
Table 5: A study of scheduler hyperparameter effect on fingerprinting
Accuracy FER
guidance scale 5 7.5 10 5 7.5 10
WOUAF 0.99031 0.99420 0.99784 0.21517 0.14531 0.16152
PALADIN 1.00000 1.00000 1.00000 0.0092 0.00600 0.00800
Table 6: A study of guidance scale hyper parameter effect on fingerprint accuracy and FER.
[Plot showing bit accuracy versus PSNR for autoencoder proposed by Lee et al. [8]] [Plot showing bit accuracy versus PSNR for autoencoder proposed by Minnen et al. [11]]
Figure 3: Plots showing bit accuracy verus PNSR when autoencoders are used to manipulate the fingerprint.
@@ 4.3 Resilience against image post-processing
One of the most common techniques to corrupt or erase fingerprint is image post-processing techniques, in that a malicious user attempts to obscure the fingerprint by employing one or more post processing methods on the generated image whilst preserving the photo-realistic nature of image. Here, we test our methodology against multiple post processing attacks such as brightness, contrast, saturation, sharpness, horizontal flip, crop and Gaussian noise and benchmark against WOUAF. Table 4(a) and 4(b) show the minimum accuracy and maximum FER seen for a range of scale values for individual image post-processing operations, PALADIN shows better accuracy and FER against WOUAF on all image post-processing employed. It is observed that accuracy drops below 100% beyond these range of post processing scale value, but key point to note is image quality is also affected which implies fingerprint corruption is possible at the cost of image quality. We also evaluate robustness against JPEG compression, from Table 4(a) and 4(b) we infer that PALADIN performs better than WOUAF when employed with JPEG compression with jpeg quality ranged from 60 to 99.
4.4 Resilience against auto encoders.
Deep learning techniques such as auto-encoders [8] [11] are used with the aim of corrupting fingerprints in the artifacts. Figure 3 shows the plot of accuracy versus PSNR for autoencoders [8] [11]. We can observe PALADIN performing significantly better than WOUAF. A key point to note is that the accuracy is dropping beyond a certain PSNR value, indicating a degradation in image quality due to significant compression strength. This also suggest that the degradation in fingerprint accuracy is possible only by reducing image quality.
5 Resilience against model hyperparameters
Accuracy FER
WOUAF 0.99420 0.01120
PALADIN 1.00000 0.00320
Table 7: A study of negative prompt effect on fingerprint accuracy and FER.
[Plot showing bit accurary for different output image sizes] [Plot showing FER for different output image sizes]
Figure 4: Plots showing accuracy and FER for varying image sizes.
In this section, model’s hyperparameters such as guidance scale, number of inference steps, scheduler, and negative prompt are varied and tested to observe its effect on fingerprint accuracy and FER.
  Table 5 shows the effect of using different scheduler and number of inference step on accuracy and FER. The weights used in these experiments were trained on Euler scheduler and the number of inference steps was set to 7.5. Table 5(a) shows PALADIN’s accuracy being more stable and better than WOUAF, Table 5(b) shows that our methodology has the least FER with a minimal increase in FER when inference step is varied.
As second part, guidance scale is varied, the model was trained with the guidance scale 7.5. Table 6 shows the affect on fingerprint accuracy and FER when the guidance scale is varied from 5 to 10, we see that both WOUAF and our methodology have little to no variation across change in guidance scale. This behaviour is partly due to guidance having its impact on denoising UNET and not on Decoder block.
As third part, negative prompts such as "blurry image", "low quality", etc. were used to manipulate the output of the Stable Diffusion model. Table 7 our methodology and WOUAF’s accuracy do not have a significant impact in terms of accuracy and FER.
As last part, different output image sizes are experimented, Figure 4 shows the impact of fingerprint accuracy and FER on different output image size. It is observed that the accuracy is maintained well in PALADIN compared to WOUAF due to the better decoding network and also because of the error correction. FER is impacted significantly for WOUAF when non-trained image sizes are selected.
@@ 6 Conclusions
Fingerprinting strategies for latent diffusion models, even with 99% attribution accuracies, cannot be deployed due to the unacceptably large absolute number of incorrect predictions involved. We introduce a framework, PALADIN, which preserves generated image quality and also achieves 100% user-attribution accuracy. This methodology shows the performance of WOUAF can be further improved by a combination of modified fingerprint decoder architecture and a loss function better equipped to handle the trade-off between generated image quality and the attribution accuracy. Next, we further enhance the model performance by incorporating the ability of bit error recovery, a concept borrowed from the literature of coding theory. Although our methodology is built atop WOUAF, it is not restricted to this one model and provides a plug-and-play layer of cyclic error correction scheme on top of similar models and is capable of improving user-attribution accuracy, along with an ability to appropriately flag non-identifiable fingerprints. In this work, we also introduce a new metric FER for measuring fingerprinting accuracy, that is more accurate and informative. To the best of our knowledge, PALADIN is the first ever strategy to achieve cent percent user attribution accuracy in text-to-image diffusion models, which makes it possible for real-world deployment.